OpenStudy (anonymous):

Simplify tan2 θ(1 – cot2 θ).

OpenStudy (lgbasallote):

\[\tan^2 \theta (1 - \cot ^2 \theta)?\]

OpenStudy (callisto):

Hmmm I think \[1+cot^2\theta=csc^2\theta\] Since \[sin^2\theta+cos^2\theta=1\]\[\frac{sin^2\theta}{sin^2\theta}+\frac{cos^2\theta}{sin^2\theta}=\frac{1}{sin^2\theta}\]\[1+cot^2\theta=csc^2\theta\] So, the first post of @Australopithecus is not correct.., I think...

OpenStudy (australopithecus):

wait nope I'm wrong lol http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (callisto):

\[tan^2 \theta (1-cot^2 \theta) = tan^2 \theta (1)- tan^2 \theta (cot^2 \theta) = tan^2 -1\]

OpenStudy (callisto):

For the last step, I missed a theta there... Sorry

OpenStudy (australopithecus):

I probably should have reviewed trig identites before attempting this

OpenStudy (callisto):

\[tan^2 \theta -1\] is just the answer for this question :|